# Leveraging data-driven weather models for improving numerical weather prediction skill through large-scale spectral nudging

### v2

@@ Abstract
Operational meteorological forecasting has long relied on physics-based numerical weather prediction (NWP) models. Recently, this landscape is facing disruption by the advent of data-driven artificial intelligence (AI)-based weather models, which offer tremendous computational performance and competitive forecasting skill. However, data-driven models for medium-range forecasting generally suffer from major limitations, including low effective resolution and a narrow range of predicted variables. This study illustrates the relative strengths and weaknesses of these competing paradigms using the GEM (Global Environmental Multiscale) and GraphCast models to represent physics-based and AI-based approaches, respectively. By analyzing global predictions from these two models against observations and analyses in both physical and spectral spaces, this study demonstrates that GraphCast-predicted large scales outperform GEM, particularly for longer lead times. Building on this insight, a hybrid NWP-AI system is proposed, wherein GEM-predicted large-scale state variables are spectrally nudged toward GraphCast predictions, while allowing GEM to freely generate fine-scale details critical for weather extremes. Results indicate that this hybrid approach is capable of leveraging the strengths of GraphCast to enhance the prediction skill of the GEM model. Importantly, trajectories of tropical cyclones are predicted with enhanced accuracy without significant changes in intensity. Furthermore, this new hybrid system ensures that meteorologists have access to a complete set of forecast variables, including those relevant for high-impact weather events.
@@ 1 Introduction
State-of-the-art numerical weather prediction (NWP) models generally include some form of a dynamical core that solves the nonlinear partial differential equations governing the resolved-scale dynamical and thermodynamical processes within the atmosphere. Solutions from the dynamical core are subsequently combined with tendencies from parameterized physics schemes, representing diabatic, frictional and subgrid-scale processes that are not explicitly accounted for by the dynamical equations. Although statistical approaches have been explored in the past — particularly for downscaling purposes (Yu et al., 2006; Busuioc et al., 2008; Cheng et al., 2014) — physics-based NWP models have long been the foundational approach for operational meteorological forecasting. The recent emergence of data-driven models inspired by artificial intelligence (AI) has, however, started to seriously challenge this well established paradigm.
@@
Keisler (2022) presented an important breakthrough demonstrating considerable potential for weather forecasting with data-driven models. This was quickly followed by Pangu-Weather from Huawei (Bi et al., 2022, 2023), GraphCast from Google DeepMind (Lam et al., 2022, 2023) and several other models (e.g., Pathak et al. 2022; Chen et al. 2023a, b). In general, these models rely on some form of deep neural network architecture. Like any other application of AI, data-driven weather models necessitate a substantial volume of high quality training data. As a result, all currently available AI-based deterministic global weather simulators are trained on the ERA5 reanalyses (Hersbach et al., 2020) from the European Centre for Medium-Range Weather Forecasts (ECMWF), which is undoubtedly the most comprehensive resource available. The neural network weights within these models are specifically trained to make predictions that closely emulate ERA5.
One important aspect of these AI-based weather emulators that set them apart from the classical NWP models is their computational performance. They can speed up the production of multi-day forecasts by orders of magnitude while using a fraction of the computational resources. Furthermore, many of these models are not only comparable to the NWP models with respect to their prediction accuracy, in terms of a number of the performance metrics, they can even surpass the skill of the Integrated Forecasting System (IFS; Lang et al. 2023) from ECMWF, which is widely accepted as the benchmark within the NWP community (Bi et al., 2023; Lam et al., 2023; Chen et al., 2023a).
@@
Despite their advantages, AI-based weather emulators also have limitations. A major weakness of most of the currently available data-driven models is considerable fine-scale smoothing (Bi et al., 2022; Lam et al., 2022). Data-driven models based on generative AI, e.g., diffusion-based GenCast model from Google DeepMind (Price et al., 2024), can address the fine-scale smoothing problem to some extent. However, as of now, such approaches generally require considerable increase in computational cost, thereby diminishing one of the principal advantages of the AI models.
@@
Improving the nominal horizontal resolution of the AI model inferences (i.e., forecasts) is also challenging. The general construction of these models requires higher-resolution training data to enable these models to make inferences at higher resolutions. Since the ERA5 reanalysis is only available at 0.25°$\degree$ resolution, this restriction means that current global AI models cannot be trained for kilometer-scale forecasting. Another common limitation of the current generation of AI models is the limited range of predicted variables. At present, these models can only be trained to emulate fields that are present in the analyses dataset. However, training to predict non-analyzed variables present in the analyses (e.g., precipitation in ERA5) would generally lead to suboptimal inference. Furthermore, adding the full suite of physics-related prognostic variables required for operational NWP, such as categorization of cloud and precipitation types, would be cost-prohibitive with current AI model architectures. Time resolution of inferences is another constraint for most of the current AI models, which make predictions at 6-hour intervals, far coarser than physics-based NWP models. Pangu-Weather (Bi et al., 2022) provides separate models for 1-, 3-, 6-, and 24-hour predictions, but the models with shorter forecasting steps suffer from larger error accumulation. In summary, despite their strengths, the aforementioned limitations make it implausible for the AI models to replace the classical NWP approach in operational weather forecasting.
@@
The ultimate goal of this study is to develop a hybrid NWP-AI system for real-time forecasting applications that combines the strengths of the physics-based NWP approach with those of the AI-based approach while addressing the inherent limitations of the AI models. The physics-based model chosen for this study is the GEM (Global Environmental Multiscale) model (Girard et al., 2014), which is employed by Environment and Climate Change Canada (ECCC) for operational weather forecasting. The AI model in this study is represented by GraphCast from Google DeepMind. Based on internal testing at ECCC, among the different freely available open-source AI models, GraphCast demonstrates the best accuracy with respect to most of the metrics.
@@
The study aims to determine whether large-scale spectral nudging of GEM predictions toward inferences from a state-of-the-art AI-based weather emulator, such as GraphCast, can improve GEM’s predictability and offer a feasible approach for developing a suitable hybrid NWP-AI system. Spectral nudging can be seen as an indirect suboptimal data assimilation method (von Storch et al., 2000). Although the concept of spectral nudging was first proposed to control spatial computational modes in a limited-area model (LAM; Waldron et al. 1996), it is primarily used to improve dynamical downscaling of coarse-resolution atmospheric model simulations over high-resolution limited area domains (Leduc and Laprise, 2009; Husain et al., 2014). Spectral nudging has also been explored for global atmospheric hindcasting (Schubert-Frisius et al., 2017).
Spectral nudging has so far been unfeasible for real-time forecasting applications due to the absence of timely and accurate reference data. The emergence of AI models, like GraphCast, has unlocked the potential for applying spectral nudging in real-time operational weather forecasting. In this context, application of spectral nudging for enhancing real-time NWP represents an important novelty proposed by this study. Although the hybridization concept is illustrated with a specific combination of AI and NWP models, the proposed strategy is expected to be broadly applicable to any other combination, provided that the AI model offers added value over the NWP model.
To achieve the most optimal fusion of GraphCast and GEM, it is essential to ascertain the relative strengths and limitations of these models through objective evaluation against observations and analyses. A detailed comparative analysis of these models is also critical for devising the most optimal spectral nudging configuration, and therefore, constitutes a significant component of this study.
This article is organized as follows. Sections 2 and 3 provide some important background information regarding the GEM and GraphCast models, respectively. Details pertaining to the relative performances of these models are documented in section 4. Section 5 covers the concept of spectral nudging, its implementation in GEM as well as the optimal nudging configuration used in this study. Detailed evaluations of the spectral nudging-based hybrid NWP-AI system are presented in 6. Finally, section 7 summarizes the main conclusions of this study and outlines potential directions for future work.
2 The NWP System
@@ 2.1 The GEM Model
The dynamical core of the GEM model solves the elastic Euler system of equations using an iteratively implicit semi-Lagrangian approach (Girard et al., 2014; Husain and Girard, 2017). The equations are first transformed from regular height coordinate to some form of a terrain-following coordinate in the vertical, denoted by $\zeta$ in the model, which for ECCC’s current operational NWP systems, is based on log-hydrostatic-pressure (Husain et al., 2021). The global domain is represented as a pair of overlapping limited-area grids in a Yin-Yang configuration (Qaddouri and Lee, 2011).
Solutions from the GEM dynamical core are combined with tendencies from parameterized physics schemes (McTaggart-Cowan et al., 2019) that represent the diabatic, frictional and subgrid-scale processes. The parameterization schemes are executed in a sequential manner. Sequential coupling (Gross et al., 2018), also referred to as split-type coupling (Husain et al., 2019), is employed to combine the parameterized physics tendencies with the resolved dynamics to obtain the complete solution for a model time step.
2.2 Global Deterministic Prediction System (GDPS)
The version of the GDPS used in this study is based on its current operational configuration (McTaggart-Cowan et al., 2019; ECCC, 2024). It has a horizontal grid spacing of approximately 15 km along with a time-step length of 450 s. In the vertical, GDPS employs 84 prognostic levels with the top approximately at 0.1 hPa.
@@ 3.1 The GraphCast Model
The GraphCast (GC) model from Google DeepMind is based on AI, and like other global AI weather models, has been trained to emulate ECMWF’s ERA5 reanalyses (Lam et al., 2022, 2023). It requires the state of the atmosphere at times $t$ and $t-\Delta t_{GC}$ to produce a forecast valid at $t+\Delta t_{GC}$, where $t$ is the initialization time and $\Delta t_{GC}$=6 hr. Although the model can be employed auto-regressively $N$ times to generate predictions valid at $t+N\Delta t_{GC}$, it is trained to minimize the mean squared error (weighted by vertical levels) for up to 12 recursive forecast steps (from 6 hr to 3 days) by considering ERA5 as the ground truth. The objective of such a training is to produce forecasts with reasonable skill for up to day 10.
@@
At its core, GC employs graph neural networks (GNNs) with an “encoder-processor-decoder” configuration (Lam et al., 2022, 2023). The input and output states of GC are represented over a 0.25°$\degree$ latitude-longitude global grid. A major advantage of GNNs is the possibility of having arbitrary range of spatial interactions. GraphCast takes advantage of this GNN feature by introducing a multi-mesh architecture based on iteratively-refined icosahedral grids (from level-0 to level-6 refinements) within the processor component. The encoder of GC maps data from the input 0.25°$\degree$ latitude-longitude grid to the internal multi-mesh of the processor, whereas the decoder brings back information from the processor multi-mesh to the output grid.
The internal multi-mesh of GC processor is spatially homogeneous. The highest resolution (level-6) mesh has a nominal resolution of 1°$\degree$ with 40,962 nodes, while the nodes of the coarse-resolution meshes are subsets of these nodes. This facilitates multi-scale graph edges that connect the highest-resolution mesh nodes to the nodes associated with the coarse-resolution meshes to achieve multi-mesh message passing. Such an approach, therefore, leads to multi-scale interactions ranging from synoptic to mesoscales.
The atmospheric state predicted by GC is represented by 6 variables defined on multiple pressure levels (temperature, $u-v$ components of wind, geopotential, specific humidity, and vertical wind speed) and 5 surface variables (2-m temperature, 10-m wind components, mean sea-level pressure, and total precipitation). In addition to the predicted variables, the input system of GC requires time-dependent forcing in the form of total incident solar radiation at the top of the atmosphere along with information regarding the local time of the day and progression of year. These time-dependent forcing terms are computed analytically. Furthermore, the input system requires some stationary input features, such as land-sea mask, surface geopotential as well as information on the local longitude and latitude of its input grid.
Pre-trained weights of GC (learned features of the GNNs) are available from Google DeepMind for two different sets of pressure levels. The version with higher vertical resolution has 37 pressure levels with the top at 1 hPa, while the coarser-resolution version has only 13 levels with the top at 50 hPa (Lam et al., 2022). It is important to note that, in addition to training on ERA5, the 13-level GC was also fine-tuned by DeepMind on initial conditions for ECMWF’s operational IFS system, from 2016 to 2021.
3.2 GraphCast-based Global Prediction System
An experimental GC-based global prediction system has been developed at ECCC. Both the 37-level and 13-level versions of GC are available. This system takes the GDPS operational analyses (Caron and Buehner, 2022) as its input. However, the GNN weights have not been subjected to any fine-tuning at ECCC. All GC predictions utilized in this study are based on this experimental system.
Based on the evaluation of GC predictions with the GDPS analyses as input, the 13-level version of GC outperforms the 37-level version for most variables, at most levels and over most regions of the world. Although a detailed analysis of this behavior is beyond the scope of this work, it is believed to be related to systematic differences in stratospheric moisture values between the initializing GDPS analyses and ERA5 reanalyses used to train GC. The previously mentioned additional fine-tuning of GC-13L by DeepMind could also have contributed to its superior performance. This study is therefore, entirely based on the 13-level version of the system and the acronym ‘GC’ is henceforth used to refer to this version.
4 Relative performances of the GDPS and GC forecasts
@@
Ten-day forecasts from GC and GDPS are initialized using the GDPS analysis every 36 h, leading to 60 cases for boreal winter 2022 (from 01 December 2021 at $0$ UTC to 27 February 2022 at 1200 UTC) and 62 cases for boreal summer 2022 (from 01 June 2022 at $0$ UTC to 31 August 2022 at 1200 UTC). The sequence of forecast integrations initialized from these dates using an atmosphere-only operational GDPS configuration is denoted GDPS-CTL.
@@
Predictions from the two systems, GDPS-CTL and GC, are first assessed for their ability to depict the full range of atmospheric circulations using variance spectra (section 4a). This analysis is aimed at ascertaining effective resolutions of the two models, which will help to fairly interpret their relative predictive skills when comparing against radiosonde observations later in section 4b.
@@ 4.1 Verification with global power spectra
Although GC employs a 0.25°$\degree$ latitude-longitude resolution grid, its predictions are subject to a smoothing effect, which increases with lead time. This is generally assumed to be associated with learning to minimize the mean square error (MSE) (e.g., Keisler (2022), Lam et al. (2023)). The smoothing effect is well indicated by departures of 120-hr GC forecasts’ global 250-hPa kinetic energy and 500-hPa temperature variance spectra from those of GDPS-CTL forecasts and the analyses (Fig. 1). For the fine scales, GraphCast’s spectra are steeper, leading to a notable variance deficiency at wavenumbers as low as 30, corresponding to lower synoptic and sub-synoptic scales ($<$1500 km). More importantly, both GDPS-CTL forecasts and analyses much more closely follow the expected $k^{-3}$ spectral slope pertaining to the 3-dimensional quasi-geostrophic flow (Charney, 1971; Skamarock et al., 2014), which indicates that GC slopes are not realistic. Hence, the effective resolution of GC, defined as the highest wave number at which the atmospheric variability is still adequately represented, is considerably lower than its nominal resolution, i.e., grid spacing. Conversely, the effective resolution of the classical NWP models is typically about 6$\sim$8 times their nominal resolution (Skamarock, 2004). With the current operational configuration, the nominal resolution for GDPS-CTL is 0.1375°$\degree$, leading to an approximate effective resolution of 90$\sim$120 km (McTaggart-Cowan et al., 2024).
@@
Figure 1: Global (a) kinetic energy power spectrum of 250-hPa horizontal wind and (b) variance spectrum of 500-hPa temperature for the GDPS analyses (black) as well as 120-hr lead time of GDPS-CTL (blue) and GC (red), averaged over 60 cases of boreal winter 2022. The dashed line indicates the -3 slope. The 5%–95% confidence interval is presented with light shadings (not discernable for the given logarithmic scale).
@@
It is important to note that the stationary effects of orography, land-sea contrast and large latitudinal climatological variations tend to be much less affected by smoothing in GC than the less predictable transient anomalies. It is therefore useful to introduce the transient spectral amplitude ratio, defined as follows:
$\gamma(n)=\left(\frac{\overline{\sigma_{x^{\prime}x^{\prime}}(n)}}{\overline{\sigma_{y^{\prime}y^{\prime}}(n)}}\right)^{1/2},$ (1)
@@
where $\sigma_{x^{\prime}x^{\prime}}$ and $\sigma_{y^{\prime}y^{\prime}}$ are real-valued non-negative functions of the total spherical wavenumber $n$, representing global spectral variances of the transient-eddy components associated with the forecast and analysis, respectively. The overbar denotes averaging over the cases for a given season. The forecast transient eddies are defined as $x^{\prime}=x-\overline{x}$, and similarly for the analysis. The optimal value of $\gamma$ is 1 for the range of scales that are adequately represented in the reference analysis. In general, for any given length scale (or wavenumber), a value of $\gamma$ larger than 1 would imply that the forecast has more variability than the analysis, whereas a value smaller than 1 would indicate variance deficiency in the forecast.
@@
Figure 2 shows the evolution of $\gamma$ with lead time for GC and GDPS-CTL forecasts against the GDPS analysis. While $\gamma$ remains close to 1 for all lead times for GDPS-CTL, indicating that its effective resolution does not change during the integration, for GC it rapidly decreases, leading to an effective resolution as low as 1000 km for 24-hr forecasts. The smoothing effect further reduces the effective resolution to 2500$\sim$2700 km before it saturates around forecast days 3$\sim$5. It is worth noting here that GC still resolves scales smaller than these limits, but only partly ($\gamma<1$). The impact of this variance-deficiency on optimal usage of GC guidance in a hybrid system will be discussed in more detail in section 5.
@@
Figure 2: Spectral (a, b) amplitude ratio $\gamma$ and (c, d) coherence $\rho$ for the global (a, c) kinetic energy spectrum of 500-hPa horizontal wind transient-eddy deviations and (b, d) variance spectrum of 500-hPa temperature transient-eddy deviations, for GDPS-CTL (blue) and for GC (red), after 24, 72, 120 and 240 hr of integration (darker to lighter shades of color). Spectra are averaged over 60 cases of boreal winter 2022. Light shadings indicate the 5%–95% confidence interval.
Figure 2 also displays the spectral coherence defined as follows:
$\rho(n)=\frac{\overline{\sigma_{x^{\prime}y^{\prime}}(n)}}{\left(\overline{\sigma_{x^{\prime}x^{\prime}}(n)}\;\overline{\sigma_{y^{\prime}y^{\prime}}(n)}\right)^{1/2}}\;,$ (2)
@@
where $\sigma_{x^{\prime}y^{\prime}}$ is a real-valued function of the total spherical wavenumber $n$, representing the transient-eddy covariance spectrum between the GDPS-CTL (or GC forecasts) and the GDPS analyses. Spectral coherence can be interpreted as a scale-dependent spatial correlation coefficient with an optimal value of 1, which would imply a perfect spatial correlation between forecast and analysis transient-eddy anomalies. Conversely, a value of 0 would imply a total de-correlation. By integrating the individual terms on the right-hand side of Eq. 2 across entire spectrum of wavenumbers, one may obtain a quantity which is analogous to the anomaly correlation, but in space.

### v3

@@ Abstract
Operational meteorological forecasting has long relied on physics-based numerical weather prediction (NWP) models. Recently, this landscape has faced disruption by the advent of data-driven artificial intelligence (AI)-based weather models, which offer tremendous computational performance and competitive forecasting accuracy. However, data-driven models for medium-range forecasting generally suffer from major limitations, including low effective resolution and a narrow range of predicted variables. This study illustrates the relative strengths and weaknesses of these competing paradigms using the physics-based GEM (Global Environmental Multiscale) and the AI-based GraphCast models. Analyses of their respective global predictions in physical and spectral space reveal that GraphCast-predicted large scales outperform GEM, particularly for longer lead times, even though fine scales predicted by GraphCast suffer from excessive smoothing. Building on this insight, a hybrid NWP-AI system is proposed, wherein temperature and horizontal wind components predicted by GEM are spectrally nudged toward GraphCast predictions at large scales, while GEM itself freely generates the fine-scale details critical for local predictability and weather extremes. This hybrid approach is capable of leveraging the strengths of GraphCast to enhance the prediction skill of the GEM model while generating a full suite of physically consistent forecast fields with a full power spectrum. Additionally, trajectories of tropical cyclones are predicted with enhanced accuracy without significant changes in intensity. Work is in progress for operationalization of this hybrid system at the Canadian Meteorological Centre.
@@ 1 Introduction
State-of-the-art physics-based NWP models include some form of a dynamical core that solves the atmospheric governing equations, and are coupled to a suite of parameterization schemes to represent diabatic, frictional, and subgrid-scale processes that are not explicitly accounted for by the dynamical equations. Although statistical alternatives have been explored in the past — particularly for downscaling purposes (Yu et al., 2006; Busuioc et al., 2008; Cheng et al., 2014) — physics-based models have long been the foundational approach for operational meteorological forecasting. The recent emergence of data-driven models inspired by artificial intelligence (AI) has, however, started to seriously challenge this well established paradigm.
@@
Keisler (2022) presented an important breakthrough demonstrating considerable potential for weather forecasting with data-driven models. This was quickly followed by Pangu-Weather from Huawei (Bi et al., 2022, 2023), GraphCast from Google DeepMind (Lam et al., 2022, 2023), and several other models (e.g., Pathak et al. 2022; Chen et al. 2023a, b). In general, these models rely on some form of deep neural network architecture. Like any other application of AI, data-driven weather models require a substantial volume of high quality training data. As a result, all currently available AI-based deterministic global weather simulators are trained on the ERA5 reanalyses (Hersbach et al., 2020) from the European Centre for Medium-Range Weather Forecasts (ECMWF), which is undoubtedly the most comprehensive resource available. The neural network weights within these data-driven models are specifically trained to make inferences (forecasts) that closely emulate ERA5.
Although training of AI-based weather emulators is computationally expensive, their exceptional computational performance during inference accelerates production times by orders of magnitude while using a fraction of the computational resources usually devoted to the physics-based models. In addition to their efficiency, standard headline scores suggest that data-driven systems generate predictions that are more accurate even than those of the Integrated Forecasting System (IFS; Lang et al. 2023) from ECMWF, the model used to generate the ERA reanalyses on which data-driven models were trained.
@@
Despite these advantages, AI-based weather emulators also have limitations. A major weakness of most of the currently available data-driven models is considerable fine-scale smoothing (Bi et al., 2022; Lam et al., 2022). Furthermore, these models are unable to accurately represent the fundamental dynamical balances in the atmosphere, leading to smoothing that is inconsistent across physically-related variables (Bonavita, 2024). It is claimed that smoothing can largely be addressed by employing a diffusion architecture (Price et al., 2024); however, such approaches currently increase computational cost considerably, thereby diminishing one of the principal advantages of the AI paradigm.
@@
Increasing the nominal horizontal resolution of AI inferences will also be challenging, as it would require higher resolution training data. Since ERA5 reanalyses are only available on a 0.25°$\degree$ grid, current global AI models cannot be trained for kilometer-scale forecasting. Current data-driven systems are also only capable of predicting a limited number of variables, a subset of those available in the training dataset. Training to predict non-analyzed variables (e.g., precipitation in ERA5) generally leads to suboptimal inference (Lam et al., 2022, 2023). Predicting the full suite of physics-related prognostic variables required of operational NWP, such as categorization of cloud and precipitation types, would also substantially increase the computational cost. Time resolution of inferences is another constraint for most of the current AI models, whose 6-hour prediction intervals are far coarser than physics-based equivalents. Pangu-Weather (Bi et al., 2022) provides separate models for 1-, 3-, 6-, and 24-hour inferences, but the models with shorter forecasting steps suffer from larger error accumulation. These limitations make it impossible for current AI-based models to completely replace operational physics-based NWP systems.
@@
The ultimate goal of this study is to develop a hybrid NWP-AI system for real-time global forecasting applications that combines the strengths of each paradigm while overcoming their individual limitations. Although the techniques developed here are generally applicable, the implementation presented in this paper is based on Environment and Climate Change Canada’s (ECCC’s) physics-based Global Environmental Multiscale (GEM) model (Girard et al., 2014) and the AI-based GraphCast model from Google DeepMind.
@@
The hybridization involves large-scale spectral nudging of GEM predictions toward GraphCast inferences. Although the concept of spectral nudging was first proposed to control spatial computational modes in a limited-area model (Waldron et al. 1996), it is primarily used to improve dynamical downscaling over high-resolution sub-domains for regional climate modelling (Leduc and Laprise, 2009; Husain et al., 2014). In general, it can be interpreted as an indirect suboptimal data assimilation method (von Storch et al., 2000). Spectral nudging has also been explored for global atmospheric hindcasting (Schubert-Frisius et al., 2017).
Spectral nudging has so far been unfeasible for real-time global forecasting applications due to the lack of a timely and accurate reference data source. However, the emergence of AI models, with their rapid inference speed and improved large-scale skill, as shown in this study, has unlocked the potential for applying spectral nudging in real-time operational global weather forecasting.
The development of the hybrid NWP-AI system presented here, begins with background information on the GEM and GraphCast models in section 2. The relative performance of these models is assessed in section 3. Section 4 introduces spectral nudging in GEM and documents an optimal nudging configuration. Detailed evaluations of the hybrid system are presented in section 5, followed by a summary of the study’s main conclusions and recommendations for future work in section 6.
2 Model descriptions
@@ 2.1 The GEM Model
The dynamical core of the GEM model solves the elastic Euler system of equations using an iteratively implicit semi-Lagrangian approach (Girard et al., 2014; Husain and Girard, 2017). The equations are first transformed from regular height coordinate to some form of a terrain-following coordinate in the vertical, denoted by $\zeta$ in the model, which for ECCC’s current operational NWP systems, is based on log-hydrostatic-pressure (Husain et al., 2021). The global domain is represented as a pair of overlapping limited-area grids in a Yin-Yang configuration (Qaddouri and Lee, 2011). Solutions from the GEM dynamical core are combined with tendencies from parameterized physics schemes (McTaggart-Cowan et al., 2019) that represent the diabatic, frictional, and subgrid-scale processes. The resulting tendencies augment the dry-dynamical solution via sequential/split coupling (Gross et al., 2018; Husain et al., 2019) to produce the complete solution for a model time step.
The GEM configuration used here is that of the current operational GDPS (Global Deterministic Prediction System; McTaggart-Cowan et al. 2019; ECCC 2024), with a horizontal grid spacing of approximately 15 km and 450 s time step. The GDPS has 84 prognostic vertical levels with the top approximately at 0.1 hPa.
@@ 2.2 The GraphCast Model
The AI-based GraphCast model from Google DeepMind has been trained to emulate ECMWF’s ERA5 reanalyses available with a horizontal grid spacing of 0.25°$\degree$ (Lam et al., 2022, 2023). It is trained to minimize the mean squared error (weighted by vertical levels) for up to 12 recursive forecast steps (from 6 hr to 3 days), and is capable of producing forecasts with reasonable accuracy up to day 10. The atmospheric state predicted by GraphCast is represented by 6 variables defined on multiple pressure levels (temperature, $u-v$ components of wind, geopotential, specific humidity, and vertical wind speed) and 5 surface variables (2-m temperature, 10-m wind components, mean sea-level pressure, and total precipitation).
@@
At its core, GraphCast employs graph neural networks (GNNs) with an “encoder-processor-decoder” configuration (Lam et al., 2022, 2023). The input and output states of GraphCast are represented over a 0.25°$\degree$ latitude-longitude global grid. A major advantage of GNNs is the possibility of having arbitrary range of spatial interactions. GraphCast takes advantage of this GNN feature by introducing a multi-mesh architecture based on iteratively-refined icosahedral grids (from level-0 to level-6 refinements) within the processor component. The encoder of GraphCast maps data from the input 0.25°$\degree$ latitude-longitude grid to the internal multi-mesh of the processor, whereas the decoder brings back information from the processor multi-mesh to the output grid.
This study employs the 13 pressure-level version of GraphCast with pre-trained weights (learned features of the GNNs) that are available from Google DeepMind. Although a 37-level version is available, only the 13-level variant has been subjected to additional fine-tuning with ECMWF’s operational analyses (2016–2021), making it more skillful than the 37-level version. In this study, GraphCast is initialized with operational GDPS analyses (Caron and Buehner, 2022) without any fine-tuning applied to its weights.
3 Relative performances of the GDPS and GraphCast forecasts
@@
Ten-day forecasts from GDPS and GraphCast are initialized every 36 h, leading to 60 cases for boreal winter 2022 (from 01 January 2022 at 1200 UTC to 31 March 2022 at $0$ UTC) and 62 cases for boreal summer 2022 (from 01 June 2022 at $0$ UTC to 31 August 2022 at 1200 UTC). The sequence of forecast integrations initialized from these dates using an atmosphere-only operational GDPS configuration is denoted GDPS-CTL.
@@
Predictions from the two systems (GDPS-CTL and GraphCast) are first assessed for their ability to depict the full range of atmospheric circulations using variance spectra (section 3a). This analysis is aimed at ascertaining effective resolutions of the two models, which will help to fairly interpret their relative predictive skills when comparing against radiosonde observations later in section 3b.
@@ 3.1 Verification with global power spectra
Although GraphCast employs a 0.25°$\degree$ latitude-longitude resolution grid, its predictions are subject to a smoothing effect that increases with lead time. This is believed to be associated with learning to minimize the mean square error (MSE) (e.g., Keisler (2022), Lam et al. (2023)). The smoothing effect is evidenced by departures of 120-hr GraphCast forecasts’ global 250-hPa kinetic energy and 500-hPa temperature variance spectra from those of GDPS-CTL forecasts, as well as GDPS and ECMWF analyses (Fig. 1). For the fine scales, GraphCast’s spectra are steeper, leading to a notable variance deficiency at wavenumbers as low as 30, corresponding to lower synoptic and sub-synoptic scales ($<$1500 km). Unlike GraphCast, both GDPS-CTL forecasts and the two analyses closely follow the expected $k^{-3}$ spectral slope for 3D quasi-geostrophic flow (Charney, 1971; Skamarock et al., 2014). The fact that the spectral variance of ECMWF analyses – on which GraphCast has been fine-tuned – closely matches that of both GDPS-CTL forecasts and GDPS analyses demonstrates that GraphCast’s fine-scale smoothing does not originate from any lack of fine-scale variance in the training data. Overall, the effective resolution, defined as the highest wave number at which the atmospheric variability is still adequately represented, is considerably lower for GraphCast than its nominal resolution, i.e., grid spacing. Conversely, the effective resolution of the classical NWP models is typically about 6$\sim$8 times their nominal resolution (Skamarock, 2004). With the current operational configuration, the nominal resolution for GDPS-CTL is 0.1375°$\degree$, leading to an approximate effective resolution of 90$\sim$120 km (McTaggart-Cowan et al., 2024).
@@
Figure 1: Global (a) kinetic energy power spectrum of 250-hPa horizontal wind and (b) variance spectrum of 500-hPa temperature for the GDPS analyses (black) and ECMWF analyses (dotted black) as well as 120-hr lead time of GDPS-CTL (blue) and GraphCast (red), averaged over 60 cases of boreal winter 2022. The dashed line indicates the -3 slope. The 5%–95% confidence interval is presented with light shadings (not discernible for the given logarithmic scale). The small- and large-scale cut-offs for the global filter (appendix B) for fair comparison at low resolution (section 3b2) are approximately 1000 km and 2500 km, respectively, as indicated with the shaded area.
@@
It is important to note that the stationary effects of orography, land-sea contrast and latitudinal climatological variations tend to be much less affected by smoothing in GraphCast than the less-predictable transient anomalies. Let $x$ and $y$ denote the forecast and analysis, respectively. The transient anomalies – which indicate deviations of individual forecasts (or analyses) from the seasonal mean state – are then defined as $x^{\prime}=x-\overline{x}$ and $y^{\prime}=y-\overline{y}$, where the overbar denotes an average over all cases within the season of interest. It is then useful to introduce the transient spectral amplitude ratio, defined as follows:
$\gamma(n)=\left(\frac{\overline{\sigma_{x^{\prime},x^{\prime}}(n)}}{\overline{\sigma_{y^{\prime},y^{\prime}}(n)}}\right)^{1/2},$ (1)
@@
where $\sigma_{x^{\prime},x^{\prime}}$ and $\sigma_{y^{\prime},y^{\prime}}$ are real-valued non-negative functions of the total spherical wavenumber $n$, representing global spectral variances of the transient-eddy components associated with the forecast and analysis, respectively. The optimal value of $\gamma$ is 1 for the range of scales that are adequately represented in the reference analysis. In general, for any given length scale (or wavenumber), a value of $\gamma$ larger than 1 would imply that the forecast has more variability than the analysis, whereas a value smaller than 1 would indicate variance deficiency in the forecast. By replacing the seasonal means with climatology and integrating the individual terms on the right-hand side of Eq. 1 across the entire spectrum of wavenumbers, one may obtain the ratio of forecast to analysis activity. Further information on forecast activity and its spectral decomposition is provided in appendix A.
@@
Figures 2a,b show the evolution of $\gamma$ with lead time for GraphCast and GDPS-CTL forecasts against the GDPS analysis. A model’s effective resolution can be obtained by identifying the smallest wavenumber at which it suffers from a considerable drop in $\gamma$. A value of $\gamma<0.9$, implies an amplitude damping of 10%, and is assumed to be the threshold for determining a model’s effective resolution in this study. While $\gamma$ remains close to 1 for all lead times for GDPS-CTL, indicating that its effective resolution does not change during the integration, for GraphCast it rapidly decreases, leading to an effective resolution as low as 1000 km for 24-hr forecasts. The smoothing effect further reduces the effective resolution to 2500$\sim$2700 km before it saturates around forecast days 3$\sim$5. It is worth noting here that GraphCast still resolves scales smaller than these limits, but only partly ($\gamma<0.9$). The impact of this variance-deficiency on optimal usage of GraphCast guidance in a hybrid system will be discussed in more detail in section 4.
@@
Figure 2: Spectral (a, b) amplitude ratio $\gamma$, and (c, d) coherence $\rho$ for the global (a, c) kinetic energy spectrum of 500-hPa horizontal wind transient-eddy deviations, and (b, d) variance spectrum of 500-hPa temperature transient-eddy deviations, for GDPS-CTL (blue) and for GraphCast (red), after 24, 72, 120, and 240 hr of integration (darker to lighter shades of color). Spectra are averaged over 60 cases of boreal winter 2022. Light shadings indicate the 5%–95% confidence interval. The dashed horizontal lines in Figs. 2a,b indicate $\gamma$=0.9. The small- and large-scale cut-offs for the global filter (appendix B) for fair comparison at low resolution (section 3b2) are approximately 1000 km and 2500 km, respectively, as indicated with the shaded area.
Figures 2c,d display the spectral coherence defined as follows:
$\rho(n)=\frac{\overline{\sigma_{x^{\prime},y^{\prime}}(n)}}{\left(\overline{\sigma_{x^{\prime},x^{\prime}}(n)}\;\overline{\sigma_{y^{\prime},y^{\prime}}(n)}\right)^{1/2}}\;,$ (2)
@@
where $\sigma_{x^{\prime},y^{\prime}}$ is a real-valued function of the total spherical wavenumber $n$, representing the transient-eddy covariance spectrum between the GDPS-CTL (or GraphCast forecasts) and the GDPS analyses. Spectral coherence can be interpreted as a scale-dependent spatial correlation coefficient with an optimal value of 1, which would imply a perfect spatial correlation between forecast and analysis transient-eddy anomalies. Conversely, a value of 0 would imply a total de-correlation. By replacing the seasonal means with climatology and integrating the individual terms on the right-hand side of Eq. 2 across the entire spectrum of wavenumbers, one may obtain a quantity which is analogous to the anomaly correlation. Further information in this regard is presented in appendix A.